OpenStudy (anonymous):

Find the area of a regular pentagon with a side of 6 ft. Give the answer to the nearest tenth.

OpenStudy (amistre64):

i believe it would be the area created by 5 isotriangles that have a base of 6

OpenStudy (amistre64):



OpenStudy (amistre64):

the length of a side would be s= 6sin(56)/sin(72) \[5*\frac{18sin(56)}{sin(72)}\to\ \frac{90sin(56)}{sin(72)}\] IF i see it right

OpenStudy (amistre64):

my 56 is bad :)

OpenStudy (amistre64):

180-70-2 110-2 = 108 108/2 = 54 change 56 to 54